0 0
Add a comment Improve this question Transcribed image text
Answer #1

a.

Physical address is 400 + 431 = 800

b.

Physical address is 400 + 900 = 1300

But segment 1 is of length 500.

logical address > segmentation length

i.e 500 > 431

So, segnemtaton fault

c.

Physical address is 2000 + 240 = 2240

logical address > segmentation length

i.e 240 > 200

So, segnemtaton fault

\color{red}Please\; upvote\;the \;solution \;if \;it \;helped.\;Thanks!?